Assume an office building is 80,000 square feet and has two tenants-A for 50,000sf and B for 30,000sf. A’s base rent is $16.00psf and B’s base rent is $18.00psf. Assume further that the building generates parking income of $150,000 and signage income of $25,000 per year. Operating expenses for the entire building are $8.00psf. What is the Net Operating Income (NOI)?
Answer #1

so as we all know that the NOI is a profitability formula which is mainly used in the real estate industry to find out how much profit we will be able to generate out of an operating building

Net Operating Income = Real Estate Revenue - Operating Expense .............................(a)

so to calculate the Real Estate Revenue we will use the Given Data

Rent rate for 50,000sf to A = $16.00psf

Thus total Rent paid by A = 16 * 50,000 = $ 800000 ..............................(1)

Similarly

Rent rate for 30,000sf to B = $18.00psf

Thus total Rent paid by B = 18 * 30,000 = 540,000 .............................(2)

Parking income generated by the building = $ 150,000 ...............................(3)

Signage income generated by building = $ 25,000 ......................................(4)

so to calculate the total revenue generated by the building we will add from equation no (1) to (4) = (1) + (2) + (3) + (4)

on putting value from the above data

= $ (800000 + 540,000 + 150,000 + 25,000) = $ 1515000 ................... (5)

Now as it is also given in the question that the operating expense in the building is $8.00psf thus total expense for 80,000 square feet. =   $8 * 80,000 = $ 640000 ..................................(6)

Now on putting the value of equation no (5) and (6) on NOI formula denoted by equation (a)

thus on putting value in equation (a)

NOI = $ 1515000 - $ 640000 = $ 875000

Thus the Net Operating Income (NOI) is $ 875000